OpenStudy (anonymous):

27x^6-9x+21x^2/3x^2 divide

OpenStudy (anonymous):

27x^6-9x+21x^2 3x(9x^5 - 3 + 7x) 9x^5 + 7x - 3 ---------------- = ------------------ = -------------- 3x^2 3x*x x

jhonyy9 (jhonyy9):

you dont missed anything ?

jhonyy9 (jhonyy9):

9x dont have exponent ?

OpenStudy (anonymous):

long division with with polynomial functions. break it into parts. 27x^6/3x^2 = 9x^4 21x^2/3x^2 = 7 -9x/3x^2 = -9x/3x^2 <---this is the remainder value since 3x^2 doesn't go into 9x so your quotient value is 9x^4+7-(9x/3x^2)

OpenStudy (anonymous):

i lied 9x/3x^2 can be expressed as 3/x...duh

OpenStudy (anonymous):

so I am a little confused going over these steps. where did the 7 go?